OpenStudy (miss.rose):

Seven less than 5 times a number is 62. What is the number? I will medal and fan if you help me!

OpenStudy (anonymous):

what's 7+62?

OpenStudy (miss.rose):

69

OpenStudy (anonymous):

now divide that by 5

OpenStudy (tygrr321):

5 X 14 = 70 70 - 7 = 63

OpenStudy (tygrr321):

ITS NOT 62 BUT THATS WAT I GOT

OpenStudy (miss.rose):

Okay. Thank you @tygrr321 I am going to ask someone else their opinion but are more than welcome to stay and see what they got. @Ghostgate

OpenStudy (tygrr321):

OK? THX

OpenStudy (miss.rose):

Seven less than 5 times a number is 62. What is the number? 7 – 5n = 62 (line 1) 7 – 7 – 5n = 62 – 7 (line 2) –5n = 55 (line 3) –5n ÷ 5 = 55 ÷ 5 (line 4) n = –11 (line 5)

OpenStudy (ghostgate):

Okay, as far as this goes: First you have to translate the question that's being asked, which you did. However, having the equation: 7 - 5n = 62 You're next step(line 2) is wrong, do you know why?

OpenStudy (phi):

the way to understand these problems is do a few of them But to help get the idea, try using numbers. For example, do you know what is 7 less than 8 ?

OpenStudy (ghostgate):

I don't know what you were doing tygrr321, but you're answer is completely off. Just to point that out there.

OpenStudy (miss.rose):

No, I don't know why my second line is incorrect. - or subtract.

OpenStudy (phi):

How about 1 less than 2 ?

OpenStudy (miss.rose):

2-1 or Subtract 1 from 2

OpenStudy (phi):

yes, that is what the English means it also works for 7 less than 8: you do 8-7

OpenStudy (phi):

now let's look at 5 times a number call the number "x" how do you write 5 times x ?

OpenStudy (miss.rose):

5*x

OpenStudy (tygrr321):

i know that ghost gate wats wrong?

OpenStudy (phi):

yes, but when it's a number times a letter we don't bother to show the multiply so just 5x

OpenStudy (tygrr321):

u dont have to rub that in

OpenStudy (phi):

in other words, 5x is how we write 5 times x in algebra

OpenStudy (miss.rose):

Ok. Next step please.

OpenStudy (phi):

now we do Seven less than 5 times a number which is 7 less than 5x we do the same as 7 less than 8 but we use 5x instead of 8

OpenStudy (miss.rose):

Ok. Next step.

OpenStudy (phi):

how do we write 7 less than 5x ?

OpenStudy (miss.rose):

7-5x Which would equal 2x, right?

OpenStudy (ghostgate):

@tygrr321 Sorry if I rubbed it in too much. Okay, acquiring 7 - 5n = 62, you do something to both sides, however it goes with what I consider the "Do the "opposite" to both sides" system or the "Opposite" system for short, in other words, you do something to both sides that is opposite to what is already going on. If you get what I mean.

OpenStudy (phi):

you did it backwards it's not totally obvious, so you have to use the same pattern as 1 less than 2 becomes 2-1 or 7 less than 8 becomes 8-7

OpenStudy (miss.rose):

@Ghostgate I get what you mean. In this case we do the opposite of it says. @phi I do not know what you mean.

OpenStudy (phi):

I mean 7 less than 5x is not 7-5x

OpenStudy (miss.rose):

@phi Ok. Then what is it?

OpenStudy (ghostgate):

Okay in corresponding to what phi mentioned, he is stating that when you come out with "7 - 5x" it wouldn't become "2x", this goes over to the correlation of "7" and "5x" aren't like terms and therefore wouldn't subtract into one another, which is why you do the "opposite" system in order to get "x" by itself.

OpenStudy (phi):

You almost have the idea. You know how do do 1 less than 2. it turns into 2 - 1 (notice the 1 is subtracted) 7 less than 5x turns into what?

OpenStudy (miss.rose):

5x-7? That is just a guess I am really not sure.

OpenStudy (phi):

yes, that is what it means If it does not make sense, play with numbers like 2 less than 3, then when you get the idea, use letters anyway so far you have 5x-7 is 62

OpenStudy (phi):

is becomes an = sign so you now have 5x-7 = 62 the next part is to "solve for x" the first step is add +7 to both sides can you do that ?

OpenStudy (ghostgate):

Okay here's a way of depicting it: "7" = 7 "less than" = subtraction, so "7 -" "5" = 5 "times a..." = multiplication so: "7 - 5*(?)" "number" = x so: "7 - 5x" "is" = "=" so: "7 - 5x =" "62" = "62" so: "7 - 5x = 62". Do you get that?

OpenStudy (miss.rose):

@phi yes. It would be 5x-7+7= 5 = 62+7= 69?

OpenStudy (ghostgate):

Now as I was stating earlier and what phi hinted towards, we now do something to both sides of the equation to get "x" by itself, which is to do the "opposite" to both sides, and as he stated, the opposite of subtraction is addition, so therefore you add 7 to both sides, giving you: 7 + 7 - 5x = 62 + 7 What do you come out with?

OpenStudy (phi):

yes except on the left side 5x-7+7 is 5x+0 (you can't ignore the x, it's really there) but 0 adding to anything does not change it. in other words 5x+0 is 5x and you have 5x= 69 last step, is divide both sides by 5

OpenStudy (phi):

\[ \frac{5}{5}x = \frac{69}{5} \] it looks like we get an ugly number.

OpenStudy (ghostgate):

So, now that you have: 5x/5 = 69/5 What does that come down to?

OpenStudy (phi):

5/5 is 1 and 1*x is x so the left side is x I assume you can divide 69 by 5

OpenStudy (miss.rose):

13.8

OpenStudy (phi):

yes, x = 13.8 a bit ugly, but 7 less than 5*13.8 should work out to be 62

OpenStudy (ghostgate):

13.8 or 13 4/5ths, correct. So: x = 13.8 or 13 4/5ths. Now, does this make sense? Or is something still puzzling?

OpenStudy (miss.rose):

I get it now. Thank you both. I might need help a little bit later but I now know who can help me with this.

OpenStudy (ghostgate):

Alright I'm glad I could help you here! Have a great day! ;) {---Ghostgate---}